270 g of glucose is added to 1 kg of water. The boiling point of this solution is
1
Expert's answer
2020-11-04T14:19:28-0500

mass of glucose = 270g

mass of water = 1kg


no. of moles of glucose ="\begin{aligned}
\dfrac{270}{180} = 1.5\ mol
\end{aligned}"


molality of glucose(m) = "\dfrac{1.5mol}{\frac{1kg}{1kg}} = 1.5mol/kg"


From, elevation in boiling point "T_b = ik_bm"

where "k_b = 0.512\ K.kg/mol"

and i (Vant Hoff's factor) = 1, since glucose is covalent


"\therefore T_b =1× 0.512 ×1.5"

"T_b = 0.768K"


"\therefore" The boiling point of the solution = 100 + 0.768 = 100.768°C